OpenStudy (anonymous):

Write the equation of an ellipse with vertices (10, 0) and (-10, 0) and co-vertices (0, 2) and (0, -2)??

OpenStudy (anonymous):

http://www.purplemath.com/modules/ellipse3.htm

OpenStudy (anonymous):

vertices are at \((-1,0)\) and \((10,0)\) so the center should be at the origin right?

OpenStudy (anonymous):

Yes, but it is (-10,0) and (10,0)

OpenStudy (anonymous):

yeah, that was a typo

OpenStudy (anonymous):

since the center is exactly half way between them it must be at \((0,0)\) so you know it is going to look like \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

and you are basically given \(a \) and \(b\) in the question do you know what they are?

OpenStudy (anonymous):

@satellite73 is A the vertices, and B the co vertices?

OpenStudy (anonymous):

\(a\) is a number, the distance between the center and the vertices in this case \(a=10\) and similarly \(b=2\)

OpenStudy (anonymous):

so then the final answer will be x^2/100+y^2/4? @satellite73

OpenStudy (anonymous):

if you write \[\frac{x^2}{100}+\frac{y^2}{4}=1\] then yes

OpenStudy (anonymous):

don't forget the \(=1\) part !!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw